bammbiii5:

A scatter plot containing the point (2, 3) has the regression equation yˆ=5x+2 . What is the residual e when x = 2? Enter your answer in the box.

Ferredoxin:

Residual= y - yˆ y=3 yˆ=5(2)+2=12 3-12=-9 -9 is the residual. A negative residual means that the LSRL overestimated the true value.